The length of a rectangle is 5m longer than its width. If the perimeter of the rectangle is 34m, find its area.
Answer by jessica43(140)

To solve this problem you need to use the given information to write two equations and then solve.
First, you know that the length of the rectangle is 5m longer than the width, which can be written as:
L = W + 5 (with W=width and L= length)
Secondly,you know that a rectangle has 4 sides with each of the opposite sides having the same length, and the total perimeter is 34m. So:
W+W+L+L = 34 or 2W + 2L = 34
Now you can plug the first equation into the second equation, replacing the L value:
2W + 2L = 34
2W + 2(W+5) = 34
2W + 2W + 10 = 34
4W + 10 = 34
4W = 24
W = 6
Now plug this into the first equation to find the length:
L = W + 5
L = 6+5
L = 11
So the length of the rectangle is 11m and the width is 6m.
Finally, use the area of a rectangle equation (A=L*W) to finish the problem:
A = L*W
A = 11*6
A = 66 meters squared
So the area of the rectangle is 66 meters squared.